OpenStudy (anonymous):

Determine whether the vectors u and v are parallel, orthogonal, or neither. u = <7, -4>, v = <-28, 16>

OpenStudy (anonymous):

\[u.v=(7)(-28)+(-4)(16)=-196-64=-260 ~\neq0\] hence u is not perpendicular to v \[u \times v=\left[\begin{matrix}i & j& k \\ 7 & -4 & 0 \\-28~&~16&0\end{matrix}\right]=i(0-0)-j(0-0)+k[(7)(16)-(-4)(-28)]\] =k(112-112) =k(0) =0 hence u is parallel to v .

OpenStudy (anonymous):

Thanks!